# Forest-Chat: Adapting Vision-Language Agents for Interactive Forest Change Analysis

### v1

@@ Abstract
The increasing availability of high-resolution satellite imagery, together with advances in deep learning, creates new opportunities for enhancing forest monitoring workflows. Two central challenges in this domain are pixel-level change detection and semantic change interpretation, particularly for complex forest dynamics. While large language models (LLMs) are increasingly adopted for data exploration, their integration with vision-language models (VLMs) for remote sensing image change interpretation (RSICI) remains underexplored, especially beyond urban environments. This paper introduces Forest-Chat, an LLM-driven agent designed for integrated forest change analysis. The proposed framework enables natural language querying and supports multiple RSICI tasks, including change detection, change captioning, object counting, deforestation percentage estimation, and change reasoning. Forest-Chat builds upon a multi-level change interpretation (MCI) vision-language backbone with LLM-based orchestration, and incorporates zero-shot change detection via AnyChange together with an interactive point-prompt interface to support fine-grained user guidance. To facilitate adaptation and evaluation in forest environments, we introduce the Forest-Change dataset, comprising bi-temporal satellite imagery, pixel-level change masks, and multi-granularity semantic change captions generated through a combination of human annotation and rule-based methods. Experimental results show that Forest-Chat achieves mIoU and BLEU-4 scores of 67.10% and 40.17% on the Forest-Change dataset, and 88.13% and 34.41% on LEVIR-MCI-Trees, a tree-focused subset of LEVIR-MCI for joint change detection and captioning. In contrast, when Forest-Chat is used in a zero-shot capacity for change detection, it achieves mIoU scores of 59.51% and 47.32% on the two datasets respectively. These findings highlight the potential of interactive, LLM-driven RSICI systems to improve accessibility, interpretability, and analytical efficiency for forest change analysis. Two new datasets and code are publicly available at https://github.com/JamesBrockUoB/ForestChat.
@@
Vision-Language models , Multi-task learning , Change interpretation , Zero-shot change detection , LLM agents
@@ 1 Introduction
Of the world’s terrestrial landmass, forests account for 31 percent of the total area, with tropical forests accounting for nearly half of this total (Canton, 2021). Forests provide habitat for up to 80% of land-based species and represented a total carbon stock of 662 gigatonnes in 2020 (Canton, 2021; Lines et al., 2022). For a multitude of species, forests provide an abundance of resources, both in terms of physical materials and wider ecosystem services such as water and nutrient recycling and climate regulation via carbon sequestration (Lines et al., 2022; Ouaknine et al., 2025; Wegler and Kuenzer, 2024). These factors contribute to the central role of forests in adapting to and mitigating climate change. However, anthropogenic pressures and the accelerating frequency of extreme weather events are detrimentally impacting both the health and integrity of global forest ecosystems (Wegler and Kuenzer, 2024). Global forest cover is in a constant state of flux due to a myriad of pressures, such as industrial logging, agricultural expansion, wildfires, extreme weather events, disease, and pests (Hansen et al., 2013).
@@
Monitoring and quantifying the types, extent, and locations of changes across vast regions is crucial for mapping and understanding forest dynamics for both policymakers and researchers (Wegler and Kuenzer, 2024). Effective monitoring of forest loss, degradation, and recovery increasingly relies on automated data collection and processing pipelines as traditional field-surveys become insufficient to meet demand (Ouaknine et al., 2025). The last decade has seen researchers become inundated with remote sensing (RS) imagery capable of providing sub-metre pixel resolution across expansive areas (Lines et al., 2022). The rich image history, spatial and temporal density, and range of freely available sensor modalities enables a better understanding of forest ecosystems as more efficient and low-cost monitoring technologies are developed and adopted (Wegler and Kuenzer, 2024; Isaienkov et al., 2020; Lines et al., 2022).
@@
As access to forest-specific datasets and benchmarks have increased, spurred on by tools such as Google Earth Engine (GEE) (Gorelick et al., 2017), a corresponding rise in deep learning (DL) and other artificial intelligence (AI) methods for forest analysis has been observed (Yun et al., 2024; Isaienkov et al., 2020). A review by Yun et al. (2024) highlights a growing publication trend in the field across multiple tasks, including but not limited to: forest resource assessment (Nguyen et al., 2024), forest ecological disturbance monitoring (Lee et al., 2023), forest parameter retrieval at individual scale (Lin et al., 2024), and tree organ-level measurement (Cai et al., 2023). A historical reliance on CNN-based architectures for the majority of data modalities and prediction tasks (e.g. classification, segmentation, object detection, and regression) has recently given way to Transformers (Vaswani et al., 2017), attention-based architectures, and hybrid architectures. Hewarathna et al. (2024) exemplify this paradigm shift by proposing a Siamese Attention U-Net model for forest ecosystem change detection.
Figure 1: Overview of the Forest-Chat system. Forest-Chat is supplied with the MCI (Liu et al., 2024) and AnyChange (Zheng et al., 2024) models (respectively designated as FC-Supervised and FC-Zero-shot when integrated into the Forest-Chat system) to operate as its eyes, and an LLM as the brain. The proposed Forest-Change and LEVIR-MCI-Trees datasets are used as a data foundation for training and adapting the MCI model for forest change analysis.
@@
The recent research trend of exploring large language models (LLMs) as assistants to human domain experts (Ge et al., 2023) has propagated out to the task of remote sensing image change interpretation (RSICI) (Li et al., 2024a; Liu et al., 2024). Specifically, Vision-Language Models (VLMs) have been applied to RS tasks to address longstanding methodological shortcomings in AI, particularly DL-based methods (Li et al., 2024c). Gaining an enhanced semantic understanding of objects and their relationships both spatially and temporally in RS scenes has largely been neglected by vision-only models (Li et al., 2024a, c). By harnessing multiple modalities and semantic reasoning capabilities, VLMs are able to establish connections between visual concepts, subsequently demonstrating improved performance, especially on out-of-distribution data (Li et al., 2024c; Mu et al., 2025). The culmination of these systems has resulted in the proposal of chatbots capable of assisting human users in RS image processing, with the capability of providing zero-shot/open-vocabulary visual understanding (Li et al., 2024c; Irvin et al., 2025; Liu et al., 2024; Du et al., 2023).
@@
Several prescient review papers have highlighted the opportunities and challenges for large foundation model-assisted remote sensing change detection (RSCD) and remote sensing image change captioning (RSICC) for augmenting the capabilities of RS domain experts (Zhang and Wang, 2024; Zou et al., 2025; Peng et al., 2025; Li et al., 2024a, b). Promising research directions include leveraging the powerful zero-shot performance of VLMs for RSCD and RSICC tasks (Peng et al., 2025; Li et al., 2024b), incorporating multi-modal data streams into AI training (Li et al., 2024a), and deeper explorations of interactive human-machine collaboration (Peng et al., 2025; Li et al., 2024a). Developing conversational agents with access to RSICI-specific tooling capable of providing interactive and comprehensive change interpretation and analysis represents a significant advancement (Liu et al., 2024; Irvin et al., 2025). Liu et al. (2022) provide valuable insights into the rapid emergence of this field by categorising the development of new spatio-temporal VLMs for RS tasks (RS-STVLMs). Models capable of dual change detection and captioning (CDC) that incorporate LLM agents capable of task orchestration and exploratory conversations are of particular interest here. The work of Change-Agent (Liu et al., 2024) specifically represents a paradigm shift in RS-STVLM development - being the first VLM capable of dual CDC to be embedded within an agent for spatio-temporal RS tasks.
While frameworks like Change-Agent demonstrate the potential of natural language interfaces for RS tasks, they often require large task-specific training datasets to align change detection outputs with semantic descriptions (e.g. TEOChat (Irvin et al., 2025)). In parallel, recent studies have questioned whether current vision–language foundation models can reliably support zero-shot change interpretation. For single-image captioning, recent works (Zhang and Wang, 2024; Zeng et al., 2024) have documented the tendency for these models to hallucinate, fail to align generated text with the image context, or produce incomplete sentences. To date, no frameworks have been proposed for purpose-built zero-shot image change captioning. Benchmarking studies have shown that generalised foundational VLMs like GPT-4V (openai et al., 2023) perform poorly on zero-shot change captioning tasks (Zhang and Wang, 2024). Descriptions are often vague, inconsistent, or fail to capture subtle or quantitative aspects of change inherent to the scene. Due to GPT-4V’s training data being predominantly based on natural images, the model’s efficacy for Earth observation tasks is limited due to weak spatial reasoning capabilities (Zhang and Wang, 2024), especially for fine-grained, quantitative-based tasks (e.g. object change counting). In contrast, recent work has demonstrated stronger potential for pixel-level change detection. Specifically, ”Segment Any Change” by Zheng et al. (2024) introduces a promptable, zero-shot semantic change detection segmentation model called AnyChange. Their work extends the foundational Segment Anything Model (SAM) (Kirillov et al., 2023) by employing a novel training-free method called bi-temporal latent matching to extract semantic similarities between two RS images. Although such methods offer training-free zero-shot capabilities, effective domain-specific change interpretation - particularly in forest applications - remains challenging due to subtle structural variations, seasonal phenology, and heterogeneous disturbance patterns, and thus still greatly benefits from purpose-built architectures and tailored datasets.
To address these challenges, this work introduces Forest-Chat (Fig. 1), a Change-Agent inspired VLM-agent system adapted to interactive forest change analysis. The Multi-level Change Interpretation (MCI) model is retained as the ‘eyes’ of the system, capable of dual CDC, whilst an LLM is employed as the ‘brain’. AnyChange is additionally incorporated into the framework to enable zero-shot and interactive point-query based change segmentation as an alternative to computationally intensive training. Finally, the system is augmented with forest-change-specific capabilities to enhance user understanding. To support the development of this system, this paper introduces the Forest-Change dataset. It is the first dataset capable of providing both pixel-level bi-temporal change masks and semantic-level change captions solely aimed at forest change analysis. The Forest-Chat agent is capable of providing a convenient and conversational interface for researchers to interactively explore forest change events whilst reducing cognitive load and manual analysis workloads. More broadly, this paradigm further opens opportunities for expert-in-the-loop analysis, allowing knowledge-driven reasoning and shared expertise to be integrated into the understanding of long-term and heterogeneous forest dynamics. Extensive experiments benchmark Forest-Chat’s supervised and zero-shot change interpretation capabilities against contemporary models on the introduced Forest-Change dataset and a LEVIR-MCI (Liu et al., 2024) dataset subset. Further ablation studies are conducted to investigate the influence of backbone model size and various multi-task learning optimisation approaches on model performance.
The contributions of this work are outlined below.
@@
    Forest-Chat is presented as an interactive forest change analysis tool that incorporates the MCI model (FC-Supervised) for pixel-level change detection and captioning (CDC), and the zero-shot AnyChange model (FC-Zero-shot) solely for pixel-level change detection. Additional tooling that further analyses the outputs of these models is provided to enable deeper forest change understanding. The LLM-driven orchestration and dialogue enables a flexible and innovative approach for exploring RS imagery. Forest-Chat can perform various tasks relating to temporal forest analysis. Extensive experiments on two disparate datasets explores several architectural components of the MCI model.
@@
    Forest-Change is proposed as a demonstration dataset for RSICI in forest applications by incorporating bi-temporal image pairs, corresponding change masks and supplementary captions that describe the degree, location, and patterns of observed deforestation.
@@
    LEVIR-MCI-Trees is additionally included as a benchmarking dataset to contrast against Forest-Change. It is a subset of LEVIR-MCI that only contains examples with captions containing tree cover changes. This larger, urban-focused dataset enables an analysis of generalisation capabilities against the smaller, scene restricted Forest-Change dataset.
@@
    Interactive point querying for the AnyChange model is made available inside the Forest-Chat interface. This extension aims to improve the user experience when utilising Forest-Chat for zero-shot change detection.
@@
    An image change captioning interface is introduced to assist with generating captions for bi-temporal image CDC datasets. Users are able to manually and automatically generate captions that can provide varied and informative descriptions pertaining to the visible changes within the RS imagery.
@@ 2 Related Work
This work concentrates on vision-language agents (VLAs) capable of multi-task learning for remote sensing image change interpretation (RSICI) tasks. Research directions into RS-VLAs are increasing in both maturity and diversity, but the areas relevant to this work include RSICI, multi-task learning (MTL), foundation and vision-language models (VLMs), and large language model (LLMs) agents.
@@ 2.1 Remote Sensing Image Change Interpretation (RSICI)
RSICI as it currently stands encompasses the tasks of RS change detection and captioning (CDC), capable of being applied across a diverse range of remote sensing data modalities and resolutions (Liu et al., 2024). RSCD is a fundamental spatio-temporal task aimed at identifying changes between multi-temporal images of a given area; either as binary predictions, or across multiple categories for deeper semantic understanding and causal attribution (Liu et al., 2025; Li et al., 2025). DL RSCD research has primarily revolved around RNN- and CNN-based models, which automatically learn hierarchical and multi-temporal features, but are constrained by local temporal and spatial receptive fields respectively. Transformers overcame this limitation by incorporating self-attention mechanisms, enabling the capture of global context and long-range dependencies that are crucial for interpreting complex scenes (Shi et al., 2024). Recently, state-space models (e.g. Mamba (Gu and Dao, 2024)) have emerged, with adaptations made for RSCD (Chen et al., 2024; Zhang et al., 2025a), which offer a promising alternative for efficient long-dependency modelling that doesn’t incur the quadratic complexity of transformers (Liu et al., 2025). Progress has been further accelerated by the widespread adoption of multi-scale feature fusion techniques, Siamese networks, and pretrained backbone networks for extracting high-level semantic features from multi-temporal RS imagery (Shi et al., 2024). Zero-shot RSCD was recently achieved by the AnyChange model (Zheng et al., 2024) which leverages the Segment Anything Model (SAM) (Kirillov et al., 2023) to provide annotation-free change analysis without task-specific fine-tuning.
@@
RSCD is fundamentally limited to explaining ”what” has changed, as compared to the deeper questions of ”why” and ”how”. The next task is to provide semantic change descriptions that bridge the gap between computer vision (CV) and natural language processing (NLP) (Shi et al., 2024). RSICC - being a combination of image captioning and change detection - demands precise visual change recognition and robust language generation capabilities to ensure the creation of correct and natural language (Liu et al., 2025; Tao et al., 2025). Contemporary methods often adopt a three-stage architecture pipeline, typically incorporating advanced CNNs, transformers, and Vision Transformers (ViTs), to extract robust semantic change representations and produce coherent textual descriptions. This three-stage paradigm was effectively established and popularised by RSICCformer (Liu et al., 2022), which coupled a CNN-based visual feature extractor with a dual-branch Siamese transformer encoder to enhance change discrimination, and a transformer caption decoder to generate coherent descriptions. The LEVIR-CC (Liu et al., 2022) benchmark dataset accompanying RSICCformer has since spurred the development of related architectures. CD4C (Li et al., 2025) integrates a supervised change-detection branch and a two-stream fusion module to adaptively handle samples with reliable or unreliable change cues. KCFI (Yang et al., 2025a) employs ViT encoders, a key-change feature perceiver, and an instruction-tuned LLM decoder jointly trained with pixel-level change detection to improve semantic grounding. This work culminates in the MCI model (Liu et al., 2024), the first model to fully unify RSCD and RSICC in a single multi-task learning framework, simultaneously generating change masks and textual descriptions, and supported by the LEVIR-MCI dataset for joint evaluation.
2.2 Multi-Task Learning (MTL)
MTL has recently been explored as a means of jointly modelling RSCD and RSICC, leveraging the complementary relationship being spatial change cues and semantic descriptions (Shi et al., 2024). Representations learned for RSCD improve localisation in captioning, while RSICC provides high-level semantic guidance that strengthens change delineation, all while reducing overall training and optimisation costs (Shi et al., 2024).
In a unified framework for RSCD and RSICC, MTL can be formalised as the joint learning of pixel-level change masks and semantic descriptions: $(M,S)=f_{CD,CC}(I)$ where $I=\{I_{t1},I_{t2},...,I_{tn}\}$ represents a temporal sequence of RS images, $M$ is a pixel-level change mask that offers granular, pixel-accurate localisation of changes, and $S$ is a high-level semantic narrative of those changes (Liu et al., 2025; Zhang and Yang, 2021).
A central challenge to MTL frameworks is ensuring that no single task dominates optimisation, since RSCD and RSICC benefit from - but also compete for - shared representations (Shi et al., 2024; Crawshaw, 2020). Existing work addresses this through architectural design choices and parameter-sharing strategies. Hard parameter sharing uses common encoders with task-specific heads, reducing overfitting risk and parameter counts. Conversely, soft parameter sharing maintains separate task-specific modules whose parameters are regularised through joint loss functions to enforce similarity, capturing shared features but increasing computational burden (Shi et al., 2024; Crawshaw, 2020).
Optimising task loss balancing is another relevant MTL research direction, with approaches evolving from inflexible manual weight assignments towards dynamic weighting strategies that adjust task contributions during training, including uncertainty-based weighting (Kendall et al., 2018) and GradNorm (Chen et al., 2018). Subsequent work tackles gradient interference explicitly by treating MTL as a multi-objective task (Sener and Koltun, 2018), or by adjusting gradient directions to promote cooperative updates (Yu et al., 2020)
2.3 Vision-Language Models (VLMs)

### v2

@@ Abstract
The increasing availability of high-resolution satellite imagery, together with advances in deep learning, creates new opportunities for forest monitoring workflows. Two central challenges in this domain are pixel-level change detection and semantic change interpretation, particularly for complex forest dynamics. While large language models (LLMs) are increasingly adopted for data exploration, their integration with vision-language models (VLMs) for remote sensing image change interpretation (RSICI) remains underexplored, especially beyond urban environments. This paper introduces Forest-Chat, an LLM-driven agent for forest change analysis, enabling natural language querying across multiple RSICI tasks, including change detection and captioning, object counting, deforestation characterisation, and change reasoning. Forest-Chat builds upon a multi-level change interpretation (MCI) vision-language backbone with LLM-based orchestration, incorporating zero-shot change detection via AnyChange and multimodal LLM-based zero-shot change captioning and refinement. To support adaptation and evaluation in forest environments, we introduce the Forest-Change dataset, comprising bi-temporal satellite imagery, pixel-level change masks, and semantic change captions generated through human annotation and rule-based methods. Forest-Chat achieves mIoU and BLEU-4 scores of 67.10% and 40.17% on Forest-Change, and 88.13% and 34.41% on LEVIR-MCI-Trees, a tree-focused subset of LEVIR-MCI. In a zero-shot capacity, it achieves 60.15% and 34.00% on Forest-Change, and 47.32% and 18.23% on LEVIR-MCI-Trees respectively. Further experiments demonstrate the value of caption refinement for injecting geographic domain knowledge into supervised captions, and the system’s limited label domain transfer onto JL1-CD-Trees. These findings demonstrate that interactive, LLM-driven systems can support accessible and interpretable forest change analysis. Datasets and code are publicly available here.
@@
Vision-Language models , Multi-task learning , Change interpretation , Zero-shot change detection and captioning , LLM agents
@@ 1 Introduction
Of the world’s terrestrial landmass, forests account for 31 percent of the total area, providing habitat for up to 80% of land-based species and representing a total carbon stock of 662 gigatonnes in 2020 (Canton, 2021; Lines et al., 2022). Forests provide an abundance of resources and ecosystem services including water and nutrient recycling and climate regulation via carbon sequestration, contributing to their central role in adapting to and mitigating climate change (Ouaknine et al., 2025). However, anthropogenic pressures and the accelerating frequency of extreme weather events - including industrial logging, agricultural expansion, and wildfires - are detrimentally impacting the health and integrity of global forest ecosystems (Wegler and Kuenzer, 2024).
@@
Monitoring and quantifying the types, extent, and locations of these changes is crucial for both policymakers and researchers (Wegler and Kuenzer, 2024). Effective monitoring increasingly relies on automated pipelines as traditional field-surveys become insufficient (Ouaknine et al., 2025), with the last decade seeing researchers inundated with remote sensing (RS) imagery capable of providing sub-meter pixel resolution across expansive areas (Lines et al., 2022). The rich image history, spatial and temporal density, and range of freely available sensor modalities enable more efficient and low-cost monitoring of forest ecosystems (Isaienkov et al., 2020).
@@
As access to forest-specific datasets and benchmarks has increased, spurred on by tools such as Google Earth Engine (GEE) (Gorelick et al., 2017), a corresponding rise in deep learning (DL) and other artificial intelligence (AI) methods for forest analysis has been observed (Isaienkov et al., 2020). A review by Yun et al. (2024) highlights a growing publication trend across multiple tasks, including forest resource assessment (Nguyen et al., 2024), ecological disturbance monitoring (Lee et al., 2023), parameter retrieval (Lin et al., 2024), and tree organ-level measurement (Cai et al., 2023). A historical reliance on CNN-based architectures has recently given way to transformers (Vaswani et al., 2017), attention-based, and hybrid architectures, exemplified by Siamese Attention U-Net models for forest change detection (Hewarathna et al., 2024).
Figure 1: Overview of the Forest-Chat system. Forest-Chat integrates multiple components: the MCI model (Liu et al., 2024) (FC-Supervised) for supervised change captioning, and the combination of AnyChange (Zheng et al., 2024) and GPT-4o (Hurst et al., 2024), which together form the FC-Zero-shot module for zero-shot change detection and captioning, as well as optional caption refinement. These components provide the system’s visual perception, while an LLM serves as the reasoning module. The proposed Forest-Change, LEVIR-MCI-Trees, and JL1-CD-Trees datasets are used as a data foundation for training and evaluating Forest-Chat for forest change analysis in a supervised and zero-shot capacity.
@@
The recent trend of exploring large language models (LLMs) as assistants to human domain experts (Ge et al., 2023) has propagated to remote sensing image change interpretation (RSICI) (Liu et al., 2024). Vision-Language Models (VLMs) have been applied to RS tasks to address longstanding methodological shortcomings in DL-based methods (Li et al., 2024c), particularly the neglect of semantic understanding of objects and their spatial and temporal relationships (Li et al., 2024a). By harnessing multiple modalities and semantic reasoning, VLMs establish connections between visual concepts and demonstrate improved performance, especially on out-of-distribution data (Mu et al., 2025). The culmination of these systems has resulted in chatbots capable of zero-shot and open-vocabulary visual understanding for RS image processing (Li et al., 2024c; Du et al., 2023).
@@
Several review papers have highlighted the opportunities and challenges for large foundation model-assisted RS change detection (RSCD) and RS image change captioning (RSICC) for augmenting domain expert capabilities (Zhang and Wang, 2024; Zou et al., 2025). Promising directions include leveraging the zero-shot performance of VLMs for RSCD and RSICC (Li et al., 2024b), incorporating multimodal data streams (Li et al., 2024a), and deeper explorations of interactive human-machine collaboration (Peng et al., 2025). Developing conversational agents with access to RSICI-specific tooling capable of providing interactive and comprehensive change interpretation represents a significant advancement (Liu et al., 2024; Irvin et al., 2025). Models capable of dual change detection and captioning (CDC) that incorporate LLM agents for task orchestration are of particular interest here - Change-Agent (Liu et al., 2024) specifically represents a paradigm shift, being the first VLM capable of dual CDC embedded within an agent for spatio-temporal RS tasks.
While frameworks like Change-Agent demonstrate the potential of natural language interfaces for RS tasks, they often require large task-specific training datasets to align change detection outputs with semantic descriptions (Irvin et al., 2025). Recent multimodal LLMs (MLLMs) capable of strong single-image RS captioning without task-specific training (Ibrahim et al., ; Zhang and Wang, 2024) offer a promising alternative, naturally motivating their application in bi-temporal change settings. However, zero-shot MLLM-based RS change captioning remains largely underexplored - where benchmarked using general prompting strategies, even large open-source MLLMs fail to capture meaningful semantic differences between bi-temporal images, producing generic or spatially inconsistent descriptions (León et al., 2026; Soni et al., 2025). This is attributed to a lack of domain-specific grounding, suggesting that further adaptation - whether through fine-tuning or domain-informed prompting - is needed to unlock the potential of MLLMs for zero-shot change captioning.
In contrast, recent work has demonstrated stronger potential for pixel-level change detection. AnyChange (Zheng et al., 2024) introduces a promptable, zero-shot semantic change detection model extending the Segment Anything Model (SAM) (Kirillov et al., 2023) by employing a training-free bi-temporal latent matching method. Although such methods offer training-free zero-shot capabilities, effective domain-specific change interpretation - particularly in forest applications - remains challenging due to subtle structural variations, seasonal phenology, and heterogeneous disturbance patterns, and thus still benefits greatly from purpose-built architectures and tailored datasets.
To address these challenges, this work introduces Forest-Chat (Fig. 1), a Change-Agent inspired VLM-agent system capable of both supervised and zero-shot interactive forest change analysis. The Multi-level Change Interpretation (MCI) model is retained as the ‘eyes’ of the system for dual supervised CDC, while an LLM serves as the ‘brain’ to orchestrate reasoning and dialogue. A zero-shot perception pipeline (FC-Zero-shot) is additionally incorporated, combining AnyChange for zero-shot and interactive point-query based change segmentation with GPT-4o for change captioning and refinement. To the best of our knowledge, Forest-Chat is the first LLM-driven agent that simultaneously integrates both supervised and zero-shot visual perception modules for RSICI. The system is augmented with analytical tools that process generated change masks to support deeper interpretation of forest change dynamics.
The paper introduces the Forest-Change dataset, the first dataset specifically designed for forest change analysis providing both bi-temporal pixel-level change masks and semantic change captions. Forest-Chat offers a conversational interface for interactive exploration of forest change events, reducing cognitive load and manual analysis effort, opening opportunities for expert-in-the-loop analysis of long-term and heterogeneous forest dynamics.
Extensive experiments evaluate Forest-Chat’s supervised and zero-shot capabilities on the Forest-Change and LEVIR-MCI-Trees datasets. Zero-shot MLLM change captioning is further benchmarked under general and style-guided prompting strategies, with the effect of domain-specific caption refinement additionally analysed. Additional studies investigate backbone model size, multi-task loss balancing and gradient conflict resolution strategies, and cross-domain transfer robustness using the JL1-CD-Trees dataset. Forest-Chat’s utility as an interactive tool for forest change analysis is further demonstrated through qualitative case studies.
The contributions of this work are outlined below:
@@
    Forest-Chat is presented as an interactive forest change analysis system integrating multiple visual perception modules, and as an empirical exploration of adapting VLM-based RS agents to forest change analysis. This is a non-trivial task given that annotated forest change data is scarce, and the ecological diversity of forest change processes — spanning abrupt clearance events to gradual degradation and reforestation — is poorly represented in existing RSICI datasets and systems. FC-Supervised provides supervised pixel-level CDC via the MCI model, while FC-Zero-shot combines AnyChange for pixel-level change detection with GPT-4o for change captioning and refinement. An LLM orchestrates reasoning and dialogue, enabling flexible exploration of temporal RS imagery. Extensive experiments across three datasets evaluate architectural components, backbone scaling, multi-task learning strategies, and cross-domain generalisation.
@@
    Zero-shot MLLM change captioning is empirically analysed as an underexplored task, benchmarking an MLLM against supervised captioning under general and style-guided prompting strategies, with the effect of domain-specific caption refinement additionally analysed.
@@
    Forest-Change is proposed as the first dedicated dataset for RSICI in forest applications, incorporating bi-temporal image pairs with corresponding change masks and captions describing the degree, location, and patterns of observed deforestation.
@@
    LEVIR-MCI-Trees is curated from LEVIR-MCI by retaining samples where tree changes are semantically present, yielding a benchmark that pairs urban-focused change detection masks with tree-aware captions. It provides a large-scale complement to Forest-Change for evaluating captioning generalisation beyond purely forest scenes.
@@
    JL1-CD-Trees is introduced as a woodland-focused subset of JL1-CD providing a rigorous out-of-distribution benchmark for cross-domain evaluation, assessing model robustness to distribution shift from sensor differences, seasonal variability, and atmospheric artifacts.
@@ 2 Related Work
This work concentrates on vision-language agents (VLAs) capable of multi-task learning for remote sensing image change interpretation (RSICI) tasks. Research directions into RS-VLAs are increasing in both maturity and diversity, but the areas relevant to this work include RSICI, foundation and vision-language models (VLMs), and large language model (LLMs) agents.
@@ 2.1 Remote Sensing Image Change Interpretation (RSICI)
RSICI commonly encompasses the tasks of RS change detection and change captioning (CDC), applicable across diverse remote sensing data modalities and resolutions (Liu et al., 2024). RSCD is a fundamental spatio-temporal task aimed at identifying changes between multi-temporal images, either as binary predictions or across multiple categories (Liu et al., 2025a; Li et al., 2025). Deep learning approaches have progressed from RNN- and CNN-based models to transformers incorporating self-attention, enabling the capture of global context and long-range dependencies crucial for complex scene interpretation (Tao et al., 2025). Progress has been further accelerated by multi-scale feature fusion, Siamese networks, and pretrained backbone networks (Shi et al., 2024). Zero-shot RSCD was recently achieved by the AnyChange model (Zheng et al., 2024), which leverages SAM (Kirillov et al., 2023) for annotation-free change analysis.
@@
RSCD is fundamentally limited to explaining what has changed, rather than why and how. RSICC addressed this by combining image captioning with change detection, demanding precise visual change recognition and robust language generation (Liu et al., 2025a; Tao et al., 2025). Contemporary methods typically adopt a three-stage architecture, incorporating CNNs, transformers, and Vision Transformers (ViTs) to extract semantic change representations and produce coherent descriptions - a paradigm established by RSICCformer (Liu et al., 2022a) and its accompanying LEVIR-CC benchmark. Subsequent work refined this paradigm - CD4C (Li et al., 2025) integrated a supervised change-detection branch to handle samples with varying change reliability, and KCFI (Yang et al., 2025a) added pixel-level change detection for improved semantic grounding - while MCI (Liu et al., 2024) was the first framework to fully unify RSCD and RSICC in a multi-task learning setting, simultaneously generating change masks and textual descriptions, supported by the LEVIR-MCI dataset. Unifying these tasks introduces the challenge of balanced optimisation, as RSCD and RSICC compete for shared representations - requiring careful consideration of loss balancing and gradient conflict resolution strategies (Shi et al., 2024; Crawshaw, 2020).
2.2 Vision-Language Models for Remote Sensing